### v8

@@ Abstract
Automatic breast cancer classification in histopathology images is crucial for precise diagnosis and treatment planning. Recently, classification approaches based on the ResNet architecture have gained popularity for significantly improving accuracy by using skip connections to mitigate vanishing gradient problems, thereby integrating low-level and high-level feature information. Nevertheless, the conventional ResNet architecture faces challenges such as data imbalance and limited interpretability, necessitating cross-domain knowledge and collaboration among medical experts. This study effectively addresses these challenges by introducing a novel method for breast cancer classification, the Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model. It integrates a pre-trained ResNet50 model with a lightweight attention mechanism, embedding an attention module in the fourth layer of ResNet50 and incorporating two fully connected layers with LeakyReLU and ReLU activation functions to enhance feature learning capabilities. The DALAResNet50 method was tested on breast cancer histopathology images from the BreakHis Database across magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. It was also compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50. The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets.
@@ III-A Transfer ResNet50
Conv1 $112\times 112112 × 112$ $7\times 77 × 7$, 64 stride=2
[…]
Pooling $1\times 11 × 1$ Average Pool, 1000-d FC.
@@ III-B3 DALAResNet50 model optimization
During the optimization phase of the DALAResNet50 model, it is adopted that focuses sequentially on pooling layers, activation layers, and nonlinear feature adjustments to enhance the model’s performance and adaptability. The related steps are as follows:
@@
    Initially, to address the issue of spatial variations in input feature maps, traditional pooling layers are replaced with adaptive average pooling. This adjustment optimizes the network’s parameter structure, enabling the model to process global contextual information more effectively and accurately. The introduction of adaptive average pooling not only simplifies the network’s complexity but also enhances the model’s ability to capture important information from different parts of the image, laying a solid foundation for subsequent feature processing.
@@
    Next, the model employs a dual activation function strategy in the activation layer, starting with the LeakyReLU activation function [10] in the initial phase. This function is chosen for its unique capability to maintain a slight yet nonzero gradient when the unit is inactive, effectively combating the problem of vanishing gradients and significantly boosting the training efficiency of deep neural networks. Following this, the ReLU activation function [11] is applied in the subsequent fully connected layer to optimize model performance further and enhance the representation of image features. The careful combination of LeakyReLU and ReLU activation functions dramatically improves the model’s ability to process and learn complex image characteristics, providing strong support for accurately classifying breast cancer images.
@@
    Finally, the introduction of element-wise multiplication operations further enhances the model’s Feature Adaptability, allowing for dynamic, nonlinear adjustments of features based on specific inputs. This step significantly increases the network’s flexibility and information utilization, enabling the model to adapt its processing flow according to the particular circumstances of the input features. This approach notably improves the model’s feature processing capability, leading to higher accuracy and efficiency in classifying breast cancer images.
Fig. 3 illustrates the flowcharts of the standard ResNet50 approach and the DALAResNet50 approach, respectively. These flowcharts highlight the newly added features in the DALAResNet50. A comprehensive comparative analysis of these differences is detailed in Table 2. The visualization of the DALAResNet50 approach is depicted in Fig. 4.
@@ IV-D Baseline selection
40X
100X
200X
400X
Fig. 6: Training and validation loss with error bars
@@ IV-E Experiment metrics
V Performance evaluation and visualization
@@
To effectively monitor the model’s performance and gain insights into the training process, it is essential to concentrate on various crucial metrics, such as training loss, validation loss, training accuracy, and validation accuracy. Visualizing these metrics offers a comprehensive view of the model’s training behavior. An in-depth analysis of these key metrics and their implications is outlined below:
@@
V-A Loss analysis
V-B Accuracy analysis
VI Summary and conclusions
This paper proposed a DALAResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale imbalanced breast cancer histopathology image datasets. The ResNet50 model incorporates the lightweight attention mechanism to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of imbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. In future research, investigating more investigations of breast multiclassification problems are planned.

### v9

@@ Abstract
Automatic breast cancer classification in histopathology images is crucial for precise diagnosis and treatment planning. Recently, classification approaches based on the ResNet architecture have gained popularity for significantly improving accuracy by using skip connections to mitigate vanishing gradient problems, thereby integrating low-level and high-level feature information. Nevertheless, the conventional ResNet architecture faces challenges such as data imbalance and limited interpretability, necessitating cross-domain knowledge and collaboration among medical experts. This study effectively addresses these challenges by introducing a novel method for breast cancer classification, the Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model. It integrates a pre-trained ResNet50 model with a lightweight attention mechanism, embedding an attention module in the fourth layer of ResNet50 and incorporating two fully connected layers with LeakyReLU and ReLU activation functions to enhance feature learning capabilities. The DALAResNet50 method was tested on breast cancer histopathology images from the BreakHis Database across magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. It was also compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50. The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets. Additionally, to address the interpretability challenges, we introduce the Dynamic Threshold Grad-CAM (DT Grad-CAM) method. DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of significant regions. This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification. The integration of DT Grad-CAM with DALAResNet50 not only ensures high classification performance but also offers enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis.
@@ II-C4 Lack of feature adaptability
II-D Grad-CAM-based Visualization Techniques in Histopathology Image Analysis
Grad-CAM (Gradient-weighted Class Activation Mapping) is a widely used technique for visualizing the regions of an image that are most relevant to a neural network’s predictions. This technique has been extensively explored in the field of histopathology image analysis to improve model interpretability and provide insights into the decision-making process of deep learning models [59].
In histopathology, accurate interpretation of tissue images is crucial for diagnosing diseases such as cancer. Traditional Grad-CAM methods generate heatmaps by computing the gradients of the target class with respect to the feature maps of the last convolutional layer, highlighting the regions of the image that contribute most to the classification decision. This allows pathologists to understand which areas of the tissue image the model focuses on, aiding in verifying the model’s predictions.
However, conventional Grad-CAM has limitations, particularly in dealing with noisy activations and less distinct regions, which can hinder the clarity and precision of the visualizations. Various enhancements have been proposed to address these challenges. For instance, Guided Grad-CAM combines Grad-CAM with guided backpropagation to produce high-resolution class-discriminative visualizations, offering finer details of the important regions [60]. Additionally, Score-CAM removes the dependency on gradients by using the target class score, improving the heatmaps’ stability and robustness[61]. “
To further enhance the interpretability and effectiveness of visual explanations in histopathology image analysis, we chose to implement the Dynamic Threshold Grad-CAM (DT Grad-CAM) method. DT Grad-CAM addresses the limitations of conventional Grad-CAM by applying adaptive thresholding using Otsu’s method, which dynamically determines the optimal threshold for highlighting significant regions. This approach reduces noise and enhances the clarity of the visualizations, making it easier to identify and interpret the critical features that contribute to the model’s predictions. By focusing on the most relevant areas and minimizing the influence of less important regions, DT Grad-CAM provides more precise and interpretable heatmaps, which are particularly beneficial in clinical settings where accurate and transparent decision-making is essential.
@@ III-A Transfer ResNet50
Conv1 $112\times 112$ $7\times 7$, 64 stride=2
Conv2_x $56\times 56$ $\begin{Bmatrix}1\times 1,&64\\ 3\times 3&64\\ 1\times 1&256\end{Bmatrix}$ x3
Conv3_x $28\times 28$ $\begin{Bmatrix}1\times 1,&128\\ 3\times 3&128\\ 1\times 1&512\end{Bmatrix}$ x4
Conv4_x $14\times 14$ $\begin{Bmatrix}1\times 1,&256\\ 3\times 3&256\\ 1\times 1&1024\end{Bmatrix}$ x6
Conv5_x $7\times 7$ $\begin{Bmatrix}1\times 1,&512\\ 3\times 3&512\\ 1\times 1&2048\end{Bmatrix}$ x3
Pooling $1\times 1$ Average Pool, 1000-d FC.
@@ III-B3 DALAResNet50 model optimization
During the optimization phase of the DALAResNet50 model, a sequential focus is placed on pooling layers, activation layers, and nonlinear feature adjustments to enhance the model’s performance and adaptability. The related steps are as follows:
@@
    Initially, to address the issue of spatial variations in input feature maps, traditional pooling layers are replaced with adaptive average pooling. This adjustment optimizes the network’s parameter structure, enabling the model to process global contextual information more effectively and accurately. The introduction of adaptive average pooling not only simplifies the network’s complexity but also enhances the model’s ability to capture important information from different parts of the image, laying a solid foundation for subsequent feature processing. Specifically, the adaptive average pooling operation is defined as follows:
    $y_{i}=\frac{1}{H_{i}\times W_{i}}\sum_{h=1}^{H_{i}}\sum_{w=1}^{W_{i}}x_{i,h,w}$ (1)
    where $y_{i}$ is the output of the adaptive average pooling layer for the $i$-th feature map, $H_{i}$ and $W_{i}$ are the height and width of the $i$-th feature map, respectively, and $x_{i,h,w}$ represents the value at position $(h,w)$ in the $i$-th feature map. This approach establishes a solid foundation for subsequent feature processing, allowing the model to effectively capture and utilize global contextual information from the input images.
@@
    Next, the model employs a dual activation function strategy in the activation layer, starting with the LeakyReLU activation function [10] in the initial phase. This function is chosen for its unique capability to maintain a slight yet nonzero gradient when the unit is inactive, effectively combating the problem of vanishing gradients and significantly boosting the training efficiency of deep neural networks. The LeakyReLU activation function is defined as follows:
    $f(x)=\begin{cases}x&\text{if }x>0\\ \alpha x&\text{otherwise}\end{cases}$ (2)
    where $f(x)$ represents the LeakyReLU function and $\alpha$ is a small constant (typically 0.01) that allows a small gradient when the input $x$ is negative. Following this, the ReLU activation function [11] is applied in the subsequent fully connected layer to optimize model performance further and enhance the representation of image features. The ReLU activation function is defined as:
    $f(x)=\max(0,x)$ (3)
    where (f (x)) is the ReLU function, which outputs (x) if (x) is positive and 0 otherwise. The careful combination of LeakyReLU and ReLU activation functions dramatically improves the model’s ability to process and learn complex image characteristics, providing strong support for accurately classifying breast cancer images. By employing LeakyReLU in the initial phase, the model effectively addresses the vanishing gradient problem, while the subsequent use of ReLU enhances the network’s nonlinearity and expressive power, thereby improving its overall performance and robustness.
@@
    Finally, the introduction of element-wise multiplication operations further enhances the model’s feature adaptability, allowing for dynamic, nonlinear adjustments of features based on specific inputs. This step significantly increases the network’s flexibility and information utilization, enabling the model to adapt its processing flow according to the particular circumstances of the input features. This approach notably improves the model’s feature processing capability, leading to higher accuracy and efficiency in classifying breast cancer images. The element-wise multiplication operation is defined as follows:
    $z_{i}=x_{i}\cdot y_{i}$ (4)
    where $z_{i}$ is the result of the element-wise multiplication of the $i$-th feature map, and $x_{i}$ and $y_{i}$ are the corresponding elements of the feature maps being multiplied. This operation allows the model to adjust the features dynamically, enhancing its ability to capture complex patterns and interactions within the data.
Fig. 3 illustrates the flowcharts of the standard ResNet50 approach and the DALAResNet50 approach, respectively. These flowcharts highlight the newly added features in the DALAResNet50. Table 2 details a comprehensive comparative analysis of these differences. Fig. 4 depicts the visualization of the DALAResNet50 approach.
@@ IV-A BreakHis Dataset
Since the BreakHis dataset is a publicly available dataset, this study does not require ethical approval. Additionally, publicly available datasets are typically subject to stringent quality control measures; therefore, this study does not separately consider image quality issues.
@@ IV-E Experiment metrics
IV-F Ablation study
In this ablation study, we compared the performance of the DALAResNet50 model and the standard ResNet50 model for benign and malignant classification at 40x, 100x, 200x, and 400x magnifications. As shown in Table 6, the DALAResNet50 model outperformed the standard ResNet50 model across all metrics, including accuracy, precision, recall, F1 score, IBA, and GMean. Additionally, we provided confusion matrices for both models on the test set, illustrated in Fig. 6. The results demonstrated that the DALAResNet50 model exhibited higher sensitivity and specificity at all magnifications, significantly reducing the number of false positives and false negatives. These findings suggest that the DALAResNet50 model surpasses the standard ResNet50 model in terms of classification accuracy and reliability.
V DT Grad-CAM
The interpretability of deep learning models is crucial, particularly in areas such as breast cancer diagnosis. Grad-CAM is a popular technique for visualizing class-specific regions of interest in images. This article introduces an enhancement called DT Grad-CAM (Dynamic Threshold Grad-CAM), which improves Grad-CAM by applying adaptive thresholds using the Otsu method, enhancing the visualizations’ clarity and relevance.
V-A DT Grad-CAM method
The DT Grad-CAM method enhances the traditional Grad-CAM approach through several steps. Each step improves the visualization of class-specific regions, making the results more interpretable and relevant for diagnostic purposes.
1. Grad-CAM Calculation
Grad-CAM uses the following formula to compute the gradient-weighted feature map:
$L_{c}^{\text{Grad-CAM}}=\text{ReLU}\left(\sum_{k}\alpha_{k}^{c}A^{k}\right)$ (5)
where:
  • $L_{c}^{\text{Grad-CAM}}$ is the Grad-CAM heatmap for class $c$.
  • $A^{k}$ is the $k$-th feature map of the convolutional layer.
  • $\alpha_{k}^{c}$ is the weight for feature map $A^{k}$, defined as the global average pooling of the gradients of class $c$ with respect to $A^{k}$:
$\alpha_{k}^{c}=\frac{1}{Z}\sum_{i}\sum{j}\frac{\partial y^{c}}{\partial A_{ij}^{k}}$ (6)
where:
  • $y^{c}$ is the score for class $c$.
  • $\frac{\partial y^{c}}{\partial A_{ij}^{k}}$ is the gradient of the score for class $c$ with respect to the feature map $A^{k}$ at location $(i,j)$.
  • $Z$ is the total number of elements in the feature map (i.e., width times height).
This step leverages the gradient information to highlight the most important features contributing to the class prediction, providing an intuitive visual explanation of the model’s decision-making process.
2. Otsu Threshold Calculation
After generating the Grad-CAM heatmap, DT Grad-CAM applies Otsu’s method to determine the optimal threshold value. Otsu’s method calculates the threshold $T$ that maximizes the between-class variance:
$\sigma_{B}^{2}(T)=\omega_{1}(T)\omega_{2}(T)[\mu_{1}(T)-\mu_{2}(T)]^{2}$ (7)
where:
  • $\sigma_{B}^{2}(T)$ is the between-class variance for threshold $T$.
  • $\omega_{1}(T)$ and $\omega_{2}(T)$ are the probabilities of foreground and background pixels at threshold $T$, respectively.
    40x
    100x
    200x
    400x
    Figure 6: Comparison of the confusion matrices of the DALAResNet50 model and the standard ResNet50 model.
    40X- Original
    40X-DTGrad-CAM
    40X- Grad-CAM
    100X- Original
    100X-DTGrad-CAM
    100X- Grad-CAM
    200X- Original
    200X-DTGrad-CAM
    200X- Grad-CAM
    400X- Original
    400X-DTGrad-CAM
    400X- Grad-CAM
    Figure 7: Comparison of Original Histopathology Images with DT Grad-CAM and Grad-CAM Visualizations at Different Magnifications
  • $\mu_{1}(T)$ and $\mu_{2}(T)$ are the means of the foreground and background pixels at threshold $T$, respectively.
By iterating through all possible thresholds $T$, the one that maximizes $\sigma_{B}^{2}(T)$ is selected as the Otsu threshold.
Using Otsu’s method allows for adaptive thresholding based on the image’s specific characteristics, ensuring that the most relevant regions are highlighted dynamically rather than relying on a fixed threshold.
3. Thresholding
The Otsu threshold is then used to binarize the Grad-CAM heatmap:
$L_{c}^{\text{Thresh}}(x,y)=\begin{cases}L_{c}^{\text{Grad-CAM}}(x,y)&\text{if }L_{c}^{\text{Grad-CAM}}(x,y)>T\\ 0&\text{otherwise}\end{cases}$ (8)
where:
  • $L_{c}^{\text{Thresh}}$ is the thresholded Grad-CAM heatmap.
  • $T$ is the optimal threshold determined by Otsu’s method.
This thresholding step ensures that only the most significant regions in the Grad-CAM heatmap are highlighted, improving the visualization’s clarity and focus by removing less relevant areas.
V-B Comparison between DT Grad-CAM and Grad-CAM Visualization
Fig. 5 illustrates a side-by-side comparison of original images, DT Grad-CAM visualizations, and traditional Grad-CAM visualizations at various magnifications (40X, 100X, 200X, and 400X). Each row shows the original image, the DT Grad-CAM visualization, and the Grad-CAM visualization, respectively. The DT Grad-CAM method consistently provides clearer and more focused representations of significant regions across all magnifications. By applying adaptive thresholding using Otsu’s method, DT Grad-CAM reduces noise and enhances the visibility of critical areas, which is particularly evident when comparing the middle columns (DT Grad-CAM) with the right columns (Grad-CAM). This adaptive approach results in sharper and more interpretable visualizations, highlighting the relevant features more distinctly and minimizing the influence of irrelevant regions. Consequently, DT Grad-CAM offers superior clarity and relevance, making it a valuable tool for improving model interpretability, especially in diagnostic applications such as breast cancer imaging.
Table 8 provides a detailed comparison between DT Grad-CAM and Grad-CAM, highlighting the key differences and advantages of DT Grad-CAM. The adaptive thresholding in DT Grad-CAM enhances the clarity and relevance of the visualizations by minimizing noise and focusing on the most significant features. The clarity of visualization is significantly improved in DT Grad-CAM by isolating the key areas of interest and reducing the influence of less relevant regions. This enhancement in clarity directly translates to better interpretability, as DT Grad-CAM offers a sharper and more meaningful highlight of critical regions, making it easier for users to understand and interpret the model’s focus areas. Noise reduction is another major advantage of DT Grad-CAM, as it effectively reduces noise in the visualization, ensuring that only the most pertinent areas are highlighted. This is particularly useful in diagnostic applications, such as breast cancer imaging, where precise and clear visualization of important regions is crucial for accurate analysis. In contrast, traditional Grad-CAM lacks these adaptive enhancements, often resulting in more noise and less precise highlighting of important regions, which can obscure the interpretation and impact the accuracy of diagnostic applications.
TABLE VIII: Comparison between DT Grad-CAM and Grad-CAM
Feature DT Grad-CAM Grad-CAM
Adaptive Thresholding Utilizes Otsu’s method to dynamically determine the optimal threshold, enhancing the clarity and relevance of the highlighted regions by minimizing noise and focusing on the most significant features. Does not employ adaptive thresholding, often leading to more noise and less precise highlighting of important regions.
Clarity of Visualization Provides clearer and more focused visualizations by isolating the key areas of interest and reducing the influence of less relevant regions. Visualizations may be more dispersed and less distinct, making it harder to interpret the significant regions accurately.
Interpretability Enhances model interpretability by offering a sharper and more meaningful highlight of critical regions, making it easier for users to understand and interpret the model’s focus areas. Provides a basic level of interpretability but can be less effective in highlighting the most relevant features clearly.
Noise Reduction Effectively reduces noise in the visualization, ensuring that only the most pertinent areas are highlighted, thus improving the overall quality of the heatmap. Higher levels of noise can obscure the interpretation, as less important regions may also be highlighted.
Relevance in Diagnostic Applications Particularly useful in diagnostic applications, such as breast cancer imaging, where precise and clear visualization of important regions is crucial for accurate analysis. Useful but may lack the precision required for detailed diagnostic purposes, potentially impacting the accuracy of interpretations.
VI Performance evaluation and visualization
@@
To effectively monitor the model’s performance and gain insights into the training process, it is essential to concentrate on various crucial metrics, such as training loss, validation loss, training accuracy, and validation accuracy. Visualizing these metrics offers a comprehensive view of the model’s training behaviour. An in-depth analysis of these key metrics and their implications is outlined below:
VI-A Loss analysis
VI-B Accuracy analysis
VII DALAResNet50’s Future Therapeutic Impact and Integration into Existing Diagnostic Processes
DALAResNet50 model promises to significantly impact the future of breast cancer treatment by enhancing diagnostic accuracy and providing clear, interpretable visualizations through the integration of DT Grad-CAM. As healthcare systems increasingly adopt artificial intelligence (AI) technologies, This model can streamline diagnostic workflows, reduce diagnostic errors, and facilitate earlier detection of breast cancer, which is crucial for improving patient outcomes.
Incorporating DALAResNet50 into existing diagnostic processes can enhance the efficiency and accuracy of histopathology evaluations. Pathologists can leverage the model’s high classification accuracy to identify cancerous tissues more confidently, thus supporting more accurate and timely diagnoses. The use of DT Grad-CAM further aids pathologists by providing visual explanations for the model’s predictions, highlighting critical regions within histopathology images, and ensuring transparency in the diagnostic process.
Moreover, the DALAResNet50 model can serve as a valuable second-opinion tool, assisting pathologists in confirming their initial diagnoses and potentially identifying cases that may require further review. This integration can lead to a more robust diagnostic process, where AI augments human expertise, reducing the cognitive load on pathologists and minimizing the risk of oversight.
As the model continues to be refined and validated with larger and more diverse datasets, its application can extend beyond breast cancer to other types of histopathological analyses. This scalability and adaptability make DALAResNet50 a versatile tool in the broader field of medical diagnostics.
In the future, integrating DALAResNet50 with electronic health record (EHR) systems can further streamline patient management by providing seamless access to diagnostic results and AI-generated insights. This integration would facilitate multidisciplinary collaboration among healthcare providers, ensuring that treatment plans are informed by the most accurate and comprehensive diagnostic information available.
Overall, the DALAResNet50 model, combined with DT Grad-CAM’s interpretability, holds great promise for transforming breast cancer diagnosis and treatment. Its integration into existing diagnostic workflows can lead to more precise, reliable, and efficient medical care, ultimately enhancing patient outcomes and advancing the field of histopathology.
@@ VII DALAResNet50’s Future Therapeutic Impact and Integration into Existing Diagnostic Processes
40X
100X
200X
400X
Fig. 9: Training and validation loss with error bars
VIII Summary and conclusions
This paper proposed a DALAResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale imbalanced breast cancer histopathology image datasets. The ResNet50 model incorporates the lightweight attention mechanism to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper’s model performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of imbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. Furthermore, this study introduces the DT Grad-CAM method to address the interpretability challenges often associated with deep learning models. DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, which provides clearer and more focused visualizations of significant regions. This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification. The integration of DT Grad-CAM with DALAResNet50 ensures high classification performance and enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis. In future research, more investigations into breast multiclassification problems are planned.